# Mora: Enabling Generalist Video Generation via A Multi-Agent Framework

### v2

@@ Abstract
Sora is the first large-scale generalist video generation model that garnered significant attention across society. Since its launch by OpenAI in February 2024, no other video generation models have paralleled Sora’s performance or its capacity to support a broad spectrum of video generation tasks. Additionally, there are only a few fully published video generation models, with the majority being closed-source. To address this gap, this paper proposes a new multi-agent framework Mora, which incorporates several advanced visual AI agents to replicate generalist video generation demonstrated by Sora. In particular, Mora can utilize multiple visual agents and successfully mimic Sora’s video generation capabilities in various tasks, such as (1) text-to-video generation, (2) text-conditional image-to-video generation, (3) extend generated videos, (4) video-to-video editing, (5) connect videos and (6) simulate digital worlds. Our extensive experimental results show that Mora achieves performance that is proximate to that of Sora in various tasks. However, there exists an obvious performance gap between our work and Sora when assessed holistically. In summary, we hope this project can guide the future trajectory of video generation through collaborative AI agents.
Figure 1: Samples for text-to-video generation of Mora. Our approach can generate high-resolution, temporally consistent videos from text prompts. The samples shown are 1024$\times$576 resolution over 12 seconds duration at 75 frames in total.
@@ 1 Introduction
Since the release of ChatGPT in November 2022, the advent of generative AI technologies has marked a significant transformation, reshaping interactions and integrating deeply into various facets of daily life and industry OpenAI (2023a, b). In the realm of visual AI, the focus has predominantly been on image generation with models such as Midjourney Midjourney (2023), Stable Diffusion Rombach et al. (2022a), and DALL-E 3 Betker et al. (2023) leading the way. In contrast, video generation has remained unexplored, especially when compared to the progress seen in language and image synthesis technologies. Although recent notable video generation models like Pika pik and Gen-2 Gen have demonstrated their capacity to produce diverse and high-quality videos, they have been limited by their ability to create longer-duration videos exceeding 10 seconds. This scenario has seen a revolutionary shift with the introduction of OpenAI’s Sora OpenAI (2024a), marking a new era in video generation capabilities.
Sora, a pioneering text-to-video generative model introduced by OpenAI in February 2024, distinguishes itself by its ability to convert text prompts into detailed videos, demonstrating remarkable potential in replicating physical world dynamics. This model transcends its predecessors by generating videos up to a minute long, closely aligning with the provided text descriptions OpenAI (2024a). Beyond mere text-to-video generation, Sora excels in a range of video tasks as a generalist model, including editing, connecting, and extending footage in ways previously unachieved. In addition, the generated content is known for its multi-view perspectives and fidelity to user instructions, positioning it uniquely amongst video generation models. As we look to the future, the implications of Sora and similar advanced video generation technologies are poised to make profound contributions across various sectors, including but not limited to filmmaking Zhu et al. (2023); Zhuang et al. (2024), robotics Kapelyukh et al. (2023); Liu et al. (2022), and healthcare Awasthi et al. (2023).
Despite its innovative contributions, Sora’s closed-source nature, similar to most video generation models, presents significant challenges for the academic community. The inaccessibility hinders researchers’ ability to replicate or extend Sora’s capabilities. There is a growing trend of attempts to reverse-engineer Sora, with some studies, such as Liu et al. (2024a), proposing potential techniques that might be employed within Sora, including diffusion transformers and spatial patch strategies Sohl-Dickstein et al. (2015); Ho et al. (2020); Bao et al. (2023); Peebles and Xie (2023). Despite these efforts, achieving the same level of performance and adaptability as Sora proves to be an immense challenge. The lack of access to comparable computational power and extensive training datasets further complicates these endeavors.
Tasks Example Sora Mora Others
Text-to-video Generation
Girdhar et al. (2023); Wang et al. (2023a); Chen et al. (2024); Ma et al. (2024a)
Text-conditional Image-to-Video Generation
Blattmann et al. (2023); pik ; Gen
Extend Generated Videos
Video-to-Video Editing
Molad et al. (2023); Liew et al. (2023); Ceylan et al. (2023)
Connect Videos
Chen et al. (2023a)
Simulate Digital Worlds
To address the limitations of current video generation models, we explore the potential of multi-agent collaboration Hong et al. (2023); Wu et al. (2023a) in accomplishing generalist video generation tasks. We introduce a multi-agent framework, referred to as Mora, that leverages various advanced large models to enable text-to-video capabilities similar to Sora. Specifically, we decompose video generation into several subtasks, with each subtask assigned to a dedicated agent: (1) enhancing prompts provided by the user, (2) generating an image from an input text prompt, (3) editing or refining images based on the enhanced conditioning provided by the text, (4) generating a video from the generated image, and (5) connecting two videos. By automatically organizing agents to loop and permute through these subtasks, Mora can complete a wide array of video generation tasks through a flexible pipeline, thereby meeting the diverse needs of users.
Intuitively, equipping the model with both a starting image and text simplifies the video generation process, as it primarily needs to extrapolate the future progression of the image. This method stands in contrast to direct end-to-end text-to-video approaches Singer et al. (2023); Ho et al. (2022a); Chen et al. (2023b); Wang et al. (2023a); Chen et al. (2024). Our multi-agent collaboration framework distinctively produces an intermediate image or video during inference, enabling the preservation of the visual diversity, style, and quality inherent in the text-to-image model (see examples in Figure 1). This process even facilitates editing capabilities. By effectively coordinating the efforts of text-to-image, image-to-image, image-to-video and video-to-video agents, Mora can adeptly conduct a broad spectrum of video generation tasks while offering superior editing flexibility and visual fidelity, rivaling the performance of established models like Sora. A detailed comparison of tasks between Mora and Sora is presented in Table 1. This comparison demonstrates that, through the collaboration of multiple agents, Mora is capable of accomplishing the video-related tasks that Sora can undertake. This comparison highlights Mora’s adaptability and proficiency in addressing a multitude of video generation challenges.
To comprehensively assess the efficacy of Mora, we use basic metrics in publicly available video generation benchmark Vbench Huang et al. (2023) and self-defined metrics for six tasks, including text-to-video generation, text-conditional image-to-video generation, extending generated videos, video-to-video editing, connecting videos, and simulating digital worlds. Notably, Mora achieves superior performance in the text-to-video generation task than existing open-sourced models, ranking second only to Sora. In the other tasks, Mora also delivers competitive results, underscoring the versatility and general capabilities of our framework.
We summarize our contributions as follows:
  • In this paper, we introduce Mora, a groundbreaking meta-programming framework crafted to enhance multi-agent collaboration. This framework stands out for its structured yet adaptable system of agents, paired with an intuitive interface for the configuration of components and task pipelines. These features position Mora as a prime instrument for pushing forward the boundaries of generalist video generation tasks.
  • Our research reveals that the quality of video generation can be notably improved by leveraging the automated cooperation of multiple agents, including text-to-image, image-to-image, image-to-video, and video-to-video agents. This collaborative process starts with generating an image from text, followed by using both the generated image and the input text to produce a video. The process concludes with further refinement, extension, connection, and editing of the video.
  • Mora stands out for its exceptional performance across six video-related tasks, surpassing existing open-sourced models. This impressive achievement underlines the effectiveness of Mora, showcasing its potential as a versatile framework tailored for generalist video generation. The comprehensive results not only affirm the capabilities of Mora but also position it as a groundbreaking tool in the realm of video generation, promising significant advancements in how video content is created and utilized.
@@ 2.1 Text-to-Video Generation
Generating videos based on textual descriptions has been long discussed. While early efforts in the field were primarily rooted in GANs Wang et al. (2020); Chu et al. (2020) and VQ-VAE Yan et al. (2021), recent breakthroughs in generative video models, driven by foundational work in transformer-based architectures and diffusion models, have advanced academic research. Auto-regressive transformers are early leveraged in video generation Wu et al. (2022); Hong et al. (2022a); Kondratyuk et al. (2023). These models are designed to generate video sequences in a frame-by-frame manner, predicting each new frame based on the previously generated frames. Parallelly, the adaptation of masked language models He et al. (2021) for visual contexts, as demonstrated by Gupta et al. (2022); Villegas et al. (2022); Yu et al. (2023); Gupta et al. (2023), underscores the versatility of transformers in video generation. The recently-proposed VideoPoet Kondratyuk et al. (2023) leverages an auto-regressive language model and can multitask on a variety of video-centric inputs and outputs.
@@
In another line, large-scale diffusion models Sohl-Dickstein et al. (2015); Ho et al. (2020) show competitive performance in video generation Ho et al. (2022a); Singer et al. (2022); Khachatryan et al. (2023); Wu et al. (2023b); Du et al. (2024). By learning to gradually denoise a sample from a normal distribution, diffusion models Sohl-Dickstein et al. (2015); Ho et al. (2020) implement an iterative refinement process for video synthesis. Initially developed for image generation Rombach et al. (2022b); Podell et al. (2023), they have been adapted and extended to handle the complexities of video data. This adaptation began with extending image generation principles to video Ho et al. (2022b, a); Singer et al. (2022), by using a 3D U-Net structure instead of conventional image diffusion U-Net. In the follow-up, latent diffusion models (LDMs) Rombach et al. (2022a) are integrated into video generation Zhou et al. (2022); Chen et al. (2023b); Wang et al. (2023a); Chen et al. (2024), showcasing enhanced capabilities to capture the nuanced dynamics of video content. For instance, Stable Video Diffusion Blattmann et al. (2023) can conduct multi-view synthesis from a single image while Emu Video Girdhar et al. (2023) uses just two diffusion models to generate higher-resolution videos.
@@
Researchers have delved into the potential of diffusion models for a variety of video manipulation tasks. Notably, Dreamix Molad et al. (2023) and MagicEdit Liew et al. (2023) have been introduced for general video editing, utilizing large-scale video-text datasets. Conversely, other models employ pre-trained models for video editing tasks in a zero-shot manner Ceylan et al. (2023); Couairon et al. (2023); Yang et al. (2023); Chai et al. (2023). SEINE Chen et al. (2023a) is specially designed for generative transition between scenes and video prediction. While these above-mentioned end-to-end models exhibit remarkable proficiency in specific areas, they encounter limitations in broadening their capabilities to encompass a wider range of video tasks, particularly those requiring varied types of inputs and the generation of longer-duration videos. The introduction of diffusion transformers Peebles and Xie (2023); Bao et al. (2023); Ma et al. (2024b) further revolutionized video generation, culminating in advanced solutions like Latte Ma et al. (2024a) and Sora OpenAI (2024a). Sora’s ability to produce minute-long videos of high visual quality that faithfully follow human instructions heralds a new era in video generation, promising unprecedented opportunities for creativity and expression in digital media.
@@ 2.2 AI Agents
Large models have enabled agents to excel across a broad spectrum of applications, showcasing their versatility and effectiveness. They have greatly advanced collaborative multi-agent structures for multimodal tasks in areas such as scientific research Tang et al. (2024), software development Hong et al. (2023); Qian et al. (2023) and society simulation Park et al. (2023). Compared to individual agents, the collaboration of multiple autonomous agents, each equipped with unique strategies and behaviors and engaged in communication with one another, can tackle more dynamic and complex tasks Guo et al. (2024). Through a cooperative agent framework known as role-playing, CAMEL Li et al. (2024) enables agents to collaborate and solve complex tasks effectively. Park et al. Park et al. (2023) designed a community of 25 generative agents capable of planning, communicating, and forming connections. Liang et al. Liang et al. (2023) have explored the use of multi-agent debates for translation and arithmetic problems, encouraging divergent thinking in large language models. Hong et al. Hong et al. (2023) introduced MetaGPT, which utilizes an assembly line paradigm to assign diverse roles to various agents. In this way, complex tasks can be broken down into subtasks, which makes it easy for many agents working together to complete. Xu et al. Xu et al. (2023) used a multi-agent collaboration strategy to simulate the academic peer review process. AutoGen Wu et al. (2023a) is a generic programming framework which can be used to implement diverse multi-agent applications across different domains, using a variety of agents and conversation patterns.
Motivated by existing works, we extend the principle of collaborative agents to complete vision tasks. By integrating and coordinating agents of various roles such as text-to-image and image-to-video agents, we manage to accomplish multiple video-related tasks in a modular and extensible approach.
@@ 3 Mora: A Multi-Agent Framework for Video Generation
As illustrated in Figure 2, Mora aims to provide a framework leveraging advanced AI agents to realize Text-to-Video generation. Sec 3.1 provides the agent definitions and example pipelines enabled by this framework. Sec 3.2 introduces the implementation detail of each agent.
3.1 Agent-based Video Generation

### v3

@@ Abstract
Text-to-video generation has made significant strides, but replicating the capabilities of advanced systems like OpenAI’s Sora remains challenging due to their closed-source nature. Existing open-source methods struggle to achieve comparable performance, often hindered by ineffective agent collaboration and inadequate training data quality. In this paper, we introduce Mora, a novel multi-agent framework that leverages existing open-source modules to replicate Sora’s functionalities. We address these fundamental limitations by proposing three key techniques: (1) multi-agent fine-tuning with a self-modulation factor to enhance inter-agent coordination, (2) a data-free training strategy that uses large models to synthesize training data, and (3) a human-in-the-loop mechanism combined with multimodal large language models for data filtering to ensure high-quality training datasets. Our comprehensive experiments on six video generation tasks demonstrate that Mora achieves performance comparable to Sora on VBench Huang et al. (2024), outperforming existing open-source methods across various tasks. Specifically, in the text-to-video generation task, Mora achieved a Video Quality score of 0.800, surpassing Sora’s 0.797 and outperforming all other baseline models across six key metrics. Additionally, in the image-to-video generation task, Mora achieved a perfect Dynamic Degree score of 1.00, demonstrating exceptional capability in enhancing motion realism and achieving higher Imaging Quality than Sora. These results highlight the potential of collaborative multi-agent systems and human-in-the-loop mechanisms in advancing text-to-video generation. Our code is available at https://github.com/lichao-sun/Mora.
11footnotetext: Equal contribution
@@ 1 Introduction
Generative AI technologies have significantly transformed various industries, with substantial advancements particularly notable in visual AI through image generation models like Midjourney (Midjourney, 2023), Stable Diffusion 3 (Esser et al., 2024), and DALL-E 3 (Betker et al., 2023). These models have demonstrated remarkable capabilities in generating high-quality images from textual descriptions. However, progress in text-to-video generation, especially for videos exceeding 10 seconds, has not kept pace. Recent developments such as Pika (pik, ) and Gen-3 (Gen, b) have shown potential but are limited to producing short video clips.
A major breakthrough occurred with OpenAI’s release of Sora in February 2024—a text-to-video model capable of generating minute-long videos that closely align with textual prompts. Sora excels in various video tasks, including editing, extending footage, offering multi-view perspectives, and adhering closely to user instructions (OpenAI, 2024a). Despite its impressive capabilities, Sora remains a closed-source system, which poses significant barriers to academic research and development. Its black-box nature hinders the community’s ability to study, replicate, and build upon its functionalities, thereby slowing progress in the field. Attempts to reverse-engineer Sora are exploring potential techniques like diffusion transformers and spatial patch strategies (Sohl-Dickstein et al., 2015; Ho et al., 2020; Bao et al., 2023; Peebles & Xie, 2023), but a large gap still exists due to the intensive computation required to train everything from scratch with a single model.
To address these challenges, we propose Mora, a novel multi-agent framework that leverages ideas from standardized operating procedures (SOPs) and employs multiple agents using existing open-source modules to replicate the complex functionalities of Sora. While SOPs and multi-agent systems have been utilized in text-based tasks (Hong et al., 2023), applying them to text-to-video generation presents unique challenges. Naive multi-agent frameworks (Yuan et al., 2024; Xie et al., 2024) often fail because agents lack effective collaboration mechanisms, leading to suboptimal performance. Moreover, existing multi-agent video generation approaches struggle to balance pipeline automation with the need for high-quality training data, which is critical for producing high-fidelity videos.
Collecting high-quality video data for training is time-consuming and computationally expensive (Chen et al., 2024a), and the scarcity of such data hampers the performance of open-source models. Furthermore, the quality of available datasets varies widely, making it challenging to train models that can match the performance of proprietary systems like Sora. To overcome these challenges, leveraging human-in-the-loop mechanisms for data filtering becomes essential. By integrating human expertise in the data synthesis process, we can ensure that the training data is of high quality, which is crucial for training effective video generation models.
In this paper, we address these fundamental limitations by introducing several key techniques in Mora. First, we develop a multi-agent fine-tuning approach with a novel self-modulation factor that enhances coordination among agents, allowing them to collaborate effectively to achieve common goals. Second, we employ a data-free training strategy that uses large models to synthesize training data, reducing reliance on large labeled datasets and enabling efficient model training without extensive data collection efforts. Third, we leverage human-in-the-loop mechanisms, in combination with multimodal large language models, for data filtering during the training data synthesis process. This approach ensures the quality of the synthesized training data, leading to improved performance of the video generation pipeline. We summarize the overall pipeline and supported tasks in Figure 1.
By integrating human-in-the-loop mechanisms with multimodal large language models for data filtering in our data-free training strategy, we significantly enhance the quality of the training data and, consequently, the generated videos. These techniques collectively improve inter-agent and agent-human collaboration, enhance the quality and diversity of the generated videos, and expand the system’s capabilities to match those of Sora. Our comprehensive experiments on six video generation tasks demonstrate that Mora achieves performance comparable to Sora on VBench (Huang et al., 2024), closely approaching Sora and outperforming existing open-source methods across various tasks. Specifically, in the text-to-video generation, Mora achieves a Video Quality score of 0.800, surpassing Sora’s 0.797, and outperforms all other baselines. In the Image-to-Video Generation task, Mora matches Sora in Video-Text Integration (0.90) and Motion Smoothness (0.98 vs. 0.99), and surpasses Sora in Imaging Quality (0.67 vs. 0.63) and Dynamic Degree (1.00 vs. 0.75). Additionally, in Video-to-Video Editing, Mora matches Sora in both Imaging Quality and Temporal Style, each scoring 0.52 and 0.24, respectively. These results demonstrate Mora’s ability to not only replicate but also enhance the functionalities of Sora, providing a promising platform for future research.
Figure 1: Illustration of SOPs to conduct video-related tasks in Mora.
We summarize our main contributions as follows:
  • We introduce Mora, a novel multi-agent framework that leverages existing open-source modules to replicate the functionalities of Sora in text-to-video generation.
  • We address fundamental challenges in agent collaboration for video generation through three key designs: (1) a self-modulated multi-agent fine-tuning approach with dynamic modulation factor for enhanced coordination, (2) human-in-the-loop control mechanisms enabling real-time adjustments, and (3) a data-free training strategy using large models to synthesize diverse training data. These techniques significantly improve inter-agent cooperation and output quality in multi-agent video generation systems.
  • We demonstrate through extensive experiments that Mora achieves performance comparable to Sora on standard benchmarks, advancing the field of text-to-video generation and providing an accessible platform for future research.
@@ 2.1 Text-to-Video Generation
Generating videos based on textual descriptions has been a long-discussed topic. While early efforts in the field were primarily rooted in GANs (Wang et al., 2020; Chu et al., 2020) and VQ-VAE (Yan et al., 2021), recent breakthroughs in generative video models, driven by foundational work in transformer-based architectures and diffusion models, have significantly advanced academic research. Auto-regressive transformers are early leveraged in video generation (Wu et al., 2022; Hong et al., 2022b; Kondratyuk et al., 2023). These models are designed to generate video sequences frame by frame, predicting each new frame based on the previously generated frames. Parallelly, the adaptation of masked language models (He et al., 2021) for visual contexts, as demonstrated by (Gupta et al., 2022; Villegas et al., 2022; Yu et al., 2023; Gupta et al., 2023), underscores the versatility of transformers in video generation. The recently proposed VideoPoet (Kondratyuk et al., 2023) leverages an auto-regressive language model and can multitask on a variety of video-centric inputs and outputs.
@@
In another line, large-scale diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020) show competitive performance in video generation (Ho et al., 2022a; Singer et al., 2022; Khachatryan et al., 2023; Wu et al., 2023a; Du et al., 2024). By learning to gradually denoise a sample from a normal distribution, diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020) implement an iterative refinement process for video synthesis. Initially developed for image generation (Rombach et al., 2022b; Podell et al., 2023), they have been adapted and extended to handle the complexities of video data. This adaptation began with extending image generation principles to video (Ho et al., 2022b; a; Singer et al., 2022), by using a 3D U-Net structure instead of conventional image diffusion U-Net. In the follow-up, latent diffusion models (LDMs) (Rombach et al., 2022a) are integrated into video generation (Zhou et al., 2022; Chen et al., 2023b; Wang et al., 2023a; Chen et al., 2024a), showcasing enhanced capabilities to capture the nuanced dynamics of video content. For instance, Stable Video Diffusion (Blattmann et al., 2023) can conduct multi-view synthesis from a single image while Emu Video (Girdhar et al., 2023) uses just two diffusion models to generate higher-resolution videos.
@@
Researchers have delved into the potential of diffusion models for a variety of video manipulation tasks. Notably, Dreamix (Molad et al., 2023) and MagicEdit (Liew et al., 2023) have been introduced for general video editing, utilizing large-scale video-text datasets. Conversely, other models employ pre-trained models for video editing tasks in a zero-shot manner (Ceylan et al., 2023; Couairon et al., 2023; Yang et al., 2023; Chai et al., 2023). SEINE (Chen et al., 2023c) is specially designed for generative transition between scenes and video prediction. The introduction of diffusion transformers (Peebles & Xie, 2023; Bao et al., 2023; Ma et al., 2024a) further revolutionized video generation, culminating in advanced solutions like Latte (Ma et al., 2024b) and Sora (OpenAI, 2024a). There is also utilization in a specific domain such as Bora (Sun et al., 2024) in biomedical scenarios. Sora’s ability to produce minute-long videos of high visual quality that faithfully follow human instructions heralds a new era in video generation.
@@ 2.2 AI Agents
Large models have enabled agents to excel across a broad spectrum of applications, showcasing their versatility and effectiveness. They have greatly advanced collaborative multi-agent structures for multimodal tasks in areas such as scientific research (Tang et al., 2024), software development (Hong et al., 2023; Qian et al., 2023) and society simulation (Park et al., 2023). Compared to individual agents, the collaboration of multiple autonomous agents, each equipped with unique strategies and behaviors and engaged in communication with one another, can tackle more dynamic and complex tasks (Guo et al., 2024). Through a cooperative agent framework known as role-playing,  Li et al. (2024b) enables agents to collaborate and solve complex tasks effectively. Park et al. (2023) designed a community of 25 generative agents capable of planning, communicating, and forming connections. Liang et al. (2023) have explored the use of multi-agent debates for translation and arithmetic problems, encouraging divergent thinking in large language models.  Hong et al. (2023) introduced MetaGPT, which utilizes an assembly line paradigm to assign diverse roles to various agents. In this way, complex tasks can be broken down into subtasks, which makes it easy for many agents working together to complete.  Xu et al. (2023) used a multi-agent collaboration strategy to simulate the academic peer review process. Besides, AutoGen (Wu et al., 2023b) is a generic programming framework that can be used to implement diverse multi-agent applications across different domains, using a variety of agents and conversation patterns. This motivates our focus on applying multi-agent frameworks on text-to-video generation tasks, enabling agents to collaborate seamlessly from project inception to completion.
@@ 3 Mora: A Multi-Agent Framework for Video Generation
Current text-to-video generation models directly generate videos from textual inputs, which prevents users from supervising key aspects of video quality, style, and other important elements in real-time. To address this limitation, we propose a novel multi-agent system coupled with a self-modulated training algorithm specifically designed for the video generation task. In the subsequent sections, we outline our approach in detail. Section 3.1 describes the problem and design of our multi-agent system and the architecture of our model. Finally, in Section 3.2 and Appendix A.7, we present our data-free multi-agent fine-tuning method.
3.1 Agent Architecture of Mora
In this section, we introduce the problem and our multi-agent video generation system. To address the complexity of generating high-quality, long-duration videos, we propose a multi-agent framework where the generative model $G$ is composed of $n$ collaborating agents $\{A_{1},A_{2},\cdots,A_{n}\}$. As shown in Figure 1, each agent specializes in a specific subtask and collaborates together within the video generation pipeline. We further introduce the definition of each agent below.
Problem Definition
Let $P\in\mathcal{P}$ denote a textual prompt from the space of all possible prompts $\mathcal{P}$, describing the desired video content. A video $V$ is represented as a sequence of frames $V=\{F_{1},F_{2},\cdots,F_{T}\}$, where each $F_{t}\in\mathbb{R}^{H\times W\times C}$ corresponds to an image at time step $t$, with height $H$, width $W$, and $C$ color channels. Our goal is to generate extended-length videos that are semantically consistent with the textual prompt while exhibiting high visual quality and temporal coherence. Formally, we aim to learn a generative model $G:\mathcal{P}\rightarrow\mathcal{V}$ that maps a textual prompt $P$ to a video $V$ in the space of all possible videos $\mathcal{V}$: $V=G(P)$. The quality of the generated video can be assessed using a set of metrics =$\mathcal{M}=\{m_{1},m_{2},...,m_{K}\}$, where each $m_{i}:\mathcal{V}\times\mathcal{P}\rightarrow\mathbb{R}$ evaluates a specific aspect of the video (e.g., visual quality, temporal consistency, semantic alignment with the prompt). Our objective is to maximize these quality metrics while ensuring diversity in the generated videos. In our multi-agent framework, $G$ is decomposed into a set of specialized agents $\{A_{1},A_{2},...,A_{N}\}$, each responsible for a specific subtask in the video generation process. These agents collaborate to produce the final video output, allowing for more granular control of the generation process.
Definition and Specialization of Agents.